# Beyond Diagonal RIS for 6G Non-Terrestrial Networks: Potentials and Challenges

### v1

@@ V-B Numerical Results and Discussion
This section presents the numerical results to reveal the effectiveness of the proposed LEO communication scheme in the efficient allocation of NOMA transmission power and phase response of the BD-RIS. The analysis is carried out by setting the transmit power to 20 dBm, the operating frequency to 3.5 GHz, and the elevation angles to $45^{\circ}$. The antenna gains are set to 10 dBi, with a path-loss exponent of 2.5. The dimension of the BD-RIS elements is $0.5\,\text{m}\times 0.5\,\text{m}$, and the distance from BD-RIS to the satellite is 500 km. The altitude of the satellite is set to 600 km, and the radius of Earth is considered to be 6371 km. The reflection coefficient magnitude is 0.9, the noise power is -90 dBm, and the Rician factor is set as 10.
The simulations were conducted based on Monte Carlo simulations, where the average results were obtained over independent channel realizations. Additionally, the simulation results were plotted considering the sum rate across different numbers of BD-RIS elements. Furthermore, the results were compared with relevant benchmark schemes, including the optimal BD-IRS scheme with orthogonal multiple access (OMA) technique and NOMA scheme with random BD-RIS phase response. Result in Fig.4 shows that as the number of BD-RIS elements increases, the sum rate of the system also increases. This is because the BD-RIS effectively directs the signal toward the specific user, resulting in an increase in channel gain. However, the proposed optimal BD-RIS scheme achieves high sum rate compared to the relevent benchmark schemes. The main reason of this is joint power allocation and phase response design of the proposed BD-RIS enabled NOMA LEO communication scheme.
@@ VI Challenges and Research Directions
Although the information provided in the previous sections well establishes that BD-RIS offers significant benefits over conventional diagonal RIS, it comes with limitations that open up crucial problems needing resolution to fully exploit its advantages. On one hand, there is very limited literature highlighting the key issues associated with the use of BD-RIS in actual wireless networks and providing new research directions. On the other hand, since the use of NTN for terrestrial users is still in its infancy, the issues related to the use of BD-RIS in such settings require extensive analysis. Therefore, in this paper, we have highlighted three key issues associated with the use of BD-RIS in an NTN-assisted wireless network.
VI-A Adaptive Channel Realization
@@
The accurate realization of channels is one of the key aspects that significantly supports BD-RIS in achieving its promised performance gains. The conventional channel estimation techniques used in conventional diagonal RIS, such as semi-passive methods (enabling conventional diagonal RIS to perform pilot transmission using low-power RF chains), can easily be replicated in BD-RIS at the expense of higher power consumption. However, purely passive techniques cannot be directly adopted due to variations in RF circuitry. When incorporating BD-RIS in a NTN wireless communication scenario, the channel estimation process becomes much more complicated as the signal propagation path significantly lengthens, and fast time-varying channels are introduced (considering the LEO satellite use case of NTN). Therefore, the channel estimation techniques used in conventional diagonal RIS cannot be directly adapted to the BD-RIS-assisted NTN scenario. Instead, innovative upgrades and/or new techniques are required that not only improve the accuracy of channel estimation but also support low power consumption, promoting sustainability.
VI-B RF Circuitry Constraints
@@
The development of BD-RIS is still in its early stages, where the mathematical models used to simulate it for various applications mostly rely on ideal assumptions, such as lossless impedance components, perfect matching, and no mutual coupling between BD-RIS elements. These assumptions directly impact the beamforming capability of BD-RIS by increasing the complexity of performing beamforming. Moreover, they also limit the implementation of the BD-RIS matrix using discrete values, while relying on simple quantization mechanisms. Therefore, it is necessary to consider these impairments when designing BD-RIS. For BD-RIS-assisted NTN applications, these RF circuitry constraints present additional challenges. Signals from NTN systems travel significant distances and face atmospheric turbulence, so relying on ideal models in two-way communication significantly impacts BD-RIS-assisted NTN communication. To fully exploit the benefits of BD-RIS in NTN-assisted scenarios, RF circuitry constraints need to be fully incorporated into mathematical models and simulation analyses using innovative methods.
VI-C Receiver Sensitivity
@@
One of the main concerns in NTN-assisted terrestrial users is the distance traveled by the wireless signal between the transmitter and the receiver. Although there is supposed to be an acceptable quality of LoS connection between a satellite and a ground user, the free space path loss and other atmospheric turbulence significantly degrade the signal quality. To fully exploit the benefits of BD-RIS in an NTN scenario, the sensitivity of BD-RIS needs to be effective enough to differentiate between noise and the NTN signal. In conventional diagonal RIS, significant attention has been given to overcoming this issue. For instance, in [15], the structure of conventional diagonal RIS is repurposed to address angular sensitivity by utilizing metallic vias to develop programmable meta-atoms that are insensitive to angular distortions. A similar setup can be adopted for BD-RIS, but when considering NTN-assisted BD-RIS wireless communication scenarios, this issue needs to be addressed with innovative designs and algorithms.
@@ VII Conclusion
This article proposed BD-RIS for 6G NTNs to further enhance control over the radio environment with a beyond diagonal phase response matrices. First, we provided the fundamentals of 6G NTNs and discussed the atmospheric layer and space layer, followed by various modes and advantages of BD-RIS technology. Then, we discussed different applications of BD-RIS in 6G NTNs. To investigate the performance of BD-RIS in NTNs, we also provided a case study on BD-RIS enabled NOMA LEO satellite communication. Finally, we studied challenges and future research directions in BD-RIS enabled NTNs and concluded our work.

### v2

@@ V-B Numerical Results and Discussion
This section presents the numerical results to reveal the effectiveness of the proposed LEO communication scheme in the efficient allocation of NOMA transmission power and PR of the BD-RIS. The analysis is carried out by setting the transmit power to 20 dBm, the operating frequency to 3.5 GHz, and the elevation angles to $45^{\circ}$. The antenna gains are set to 10 dBi, with a path-loss exponent of 2.5. The dimension of the BD-RIS PREs is $0.5\,\text{m}\times 0.5\,\text{m}$, and the distance from BD-RIS to the satellite is 500 km. The altitude of the satellite is set to 600 km, and the radius of Earth is considered to be 6371 km. The reflection coefficient magnitude is 0.9, the noise power is -90 dBm, and the Rician factor is set as 10. The simulations were conducted based on Monte Carlo simulations, where the average results were obtained over independent channel realizations. The simulation results were plotted considering the sum rate across different numbers of BD-RIS PREs. Furthermore, the results were compared with those CD-RIS aided NOMA LEO satellite communications.
First, we investigate the impact of LEO transmit power on the spectral efficiency of the system with the number of PREs is 80, as depicted in Fig. Fig.4 (b). It can be seen that the spectral efficiency of both NOMA BD-RIS and NOMA CD-RIS systems increases as the transmit power of LEO increases. We can also observe that NOMA BD-RIS aided LEO communications achieve significantly high spectral efficiency compared to its counterpart NOMA CD-RIS LEO communications. For instance, using the same system parameters when LEO is operating at 15 dB transmit power, the proposed NOMA BD-RIS provides 14.56 bps/Hz while NOMA-CD-RIS can only achieve 5 bps/Hz. This is because NOMA BD-RIS offers enhanced signal control by adopting a full scattering matrix, resulting in improved signal gain.
In addition, it is interesting to see the role of PREs on the system performance in both BD-RIS and CD-RIS systems, where the transmit power of LEO is set as 10 dB. The result in Fig.4 (c) shows that as the number of BD-RIS PREs increases, the spectral efficiency of both systems also increases. This is because more PREs can effectively direct the signal towards the target users, increasing channel gain. However, the proposed NOMA BD-RIS LEO communications outperforms the NOMA CD-RIS LEO communications. The main reason for this is the advanced control of the BD-RIS system over signal manipulation compared to the CD-RIS system.
@@ VI-A Challenges
Although the information provided in the previous sections well establishes that BD-RIS offers significant benefits over CD-RIS, it comes with limitations that open up crucial problems needing resolution to fully exploit its advantages. On one hand, there is very limited literature highlighting the key issues associated with the use of BD-RIS in actual wireless networks and providing new research directions. On the other hand, since the use of NTN for terrestrial users is still in its infancy, the issues related to the use of BD-RIS in such settings require extensive analysis. Therefore, in this paper, we have highlighted three key issues associated with the use of BD-RIS in an NTN-assisted wireless network.
VI-A1 Adaptive Channel Realization
@@
The accurate realization of channels is one of the key aspects that significantly supports BD-RIS in achieving its promised performance gains. The conventional channel estimation techniques used in CD-RIS, such as semi-passive methods (enabling CD-RIS to perform pilot transmission using low-power RF chains), can easily be replicated in BD-RIS at the expense of higher power consumption. However, purely passive techniques cannot be directly adopted due to variations in RF circuitry. When incorporating BD-RIS in a NTN wireless communication scenario, the channel estimation process becomes much more complicated as the signal propagation path significantly lengthens, and fast time-varying channels are introduced (considering the LEO satellite use case of NTN). Therefore, the channel estimation techniques used in CD-RIS cannot be directly adapted to the BD-RIS aided NTN scenario. Instead, innovative upgrades and/or new techniques are required that not only improve the accuracy of channel estimation but also support low power consumption, promoting sustainability.
VI-A2 RF Circuitry Constraints
@@
The development of BD-RIS is still in its early stages, where the mathematical models used to simulate it for various applications mostly rely on ideal assumptions, such as lossless impedance components, perfect matching, and no mutual coupling between BD-RIS PREs. These assumptions directly impact the BF capability of BD-RIS by increasing the complexity of performing BF. Moreover, they also limit the implementation of the BD-RIS matrix using discrete values, while relying on simple quantization mechanisms. Therefore, it is necessary to consider these impairments when designing BD-RIS. For BD-RIS aided NTN applications, these RF circuitry constraints present additional challenges. Signals from NTN systems travel significant distances and face atmospheric turbulence, so relying on ideal models in two-way communication significantly impacts BD-RIS aided NTN communication. To fully exploit the benefits of BD-RIS in NTN-assisted scenarios, RF circuitry constraints need to be fully incorporated into mathematical models and simulation analyses using innovative methods.
VI-A3 Receiver Sensitivity
@@
One of the main concerns in NTN-assisted terrestrial users is the distance traveled by the wireless signal between the transmitter and the receiver. Although there is supposed to be an acceptable quality of LoS connection between a satellite and a ground user, the free space path loss and other atmospheric turbulence significantly degrade the signal quality. To fully exploit the benefits of BD-RIS in an NTN scenario, the sensitivity of BD-RIS needs to be effective enough to differentiate between noise and the NTN signal. In CD-RIS, significant attention has been given to overcoming this issue. For instance, in [15], the structure of CD-RIS is repurposed to address angular sensitivity by utilizing metallic vias to develop programmable meta-atoms that are insensitive to angular distortions. A similar setup can be adopted for BD-RIS, but when considering NTN-assisted BD-RIS wireless communication scenarios, this issue needs to be addressed with innovative designs and algorithms.
VI-B Research Directions
VI-B1 Artificial Intelligence/Machine Learning
On one hand, as elucidated in the preceding subsections, the three predominant challenges—namely adaptive channel realization, constraints inherent to RF circuitry, and receiver sensitivity—impose significant limitations on the comprehensive adaptation of BD-RIS within the context of NTN. On the other hand, forthcoming inquiries into BD-RIS-enhanced NTN ought to concentrate on mitigating these challenges through the integration of cutting-edge methodologies. Artificial Intelligence and Machine Learning can serve a pivotal function in the optimization of BD-RIS by facilitating real-time adjustments of signal reflections, adjusting RF circuitry losses, forecasting channel fluctuations, and implementing self-optimization based on feedback mechanisms, thereby augmenting performance in dynamic operational environments.
VI-B2 Physical Layer Security
The advancement of physical layer security algorithms is paramount for the protection of sensitive communications within NTN, wherein the inherently broadcast nature of satellite-based systems renders them susceptible to unauthorized access and signal manipulation. By fortifying these algorithms, BD-RIS can enhance the security of communications, refine real-time signal reflection adjustments, and adapt to changing network conditions, thereby ensuring more reliable and efficient performance in prospective NTN scenarios.
VI-B3 Joint Communications and Sensing
Joint communication and sensing are promising research frontiers in NTNs due to the dual benefit of using a single platform for monitoring and data transmission. However, accurate signal management and the possibility of communication and sensor signal interference make this integration difficult, especially in dynamic and resource-constrained environments like space. BD-RIS improves signal directionality and intensity, enabling precise beamforming that adapts to changing conditions, which protects communication and sensor signals. The joint communications and sensing applications in NTNs with BD-RIS hold significant potential and warrant further exploration.
@@ VII Conclusion
This article proposed BD-RIS for 6G NTNs to further enhance control over the radio environment with a beyond diagonal PRMs. First, we provided the fundamentals of 6G NTNs and discussed the atmospheric layer and space layer, followed by various hardware architectures, modes and advantages of BD-RIS technology. Then, we discussed different applications of BD-RIS in 6G NTNs. To investigate the performance of BD-RIS in NTNs, we also provided a case study on BD-RIS enabled NOMA LEO satellite communication and compared its performance with NOMA CD-RIS LEO communications. Finally, we studied challenges and future research directions in BD-RIS enabled NTNs and concluded our work.
Biographies
Wali Ullah Khan [M] (waliullah.khan@uni.lu) received a Ph.D. degree in information and communication engineering from Shandong University, Qingdao, China, in 2020. He is currently working with the SIGCOM Research Group, SnT, University of Luxembourg.
Asad Mahmood [S] (asad.mahmood@uni.lu) received his Master degrees in Electrical Engineering from COMSATS University Islamabad, Wah Campus, Pakistan. He is currently pursuing the Ph.D. degree with the Interdisciplinary Centre for Security, Reliability, and Trust (SnT), University of Luxembourg.
Muhammad Ali Jamshed (muhammadali.jamshed@glasgow.ac.uk) received a Ph.D. degree from the University of Surrey, Guildford, U.K, in 2021. He is currently working with the James Watt School of Engineering, University of Glasgow, UK.
Eva Lagunas [SM] (eva.lagunas@uni.lu) received a Ph.D. degree in telecommunications engineering from the Polytechnic University of Catalonia (UPC), Barcelona, Spain, in 2014. She currently holds a research scientist position in the SIGCOM Research Group, SnT, University of Luxembourg.
Manzoor Ahmed (manzoor.achakzai@gmail.com) received his Ph.D. degree from Beijing University of Posts and Telecommunications, Beijing China, in 2015. He is currently a professor in the School of Computer and Information Science and also with the Institute for AI Industrial Technology Research, Hubei Engineering University, Xiaogan City, 432000, China
Symeon Chatzinotas [F] (symeon.chatzinotas@uni.lu) received Ph.D. degrees in electronic engineering from the University of Surrey, Guildford, United Kingdom, in 2009. He is currently a full professor or Chief Scientist I and the co-head of the SIGCOM Research Group, SnT, University of Luxembourg.